OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 83. n + (n + 2) = 83 n − (n + 1) = 83 n + (n + 1) = 83 n + n = 83

OpenStudy (anonymous):

n+n+2=83

OpenStudy (anonymous):

n + n+1 (this implies consecutive integers) So 2n + 1 = 83 and then we can say that 2n = 82 (subtracting 1 from both sides) So therefore n = 41, and as such n + 1 = 42 41 + 42 = 83 I think you will agree

OpenStudy (anonymous):

3rd one